Assume that. Fn converges to f uniformly on S if each fn is continuous at a point c of S .show that the limit function f us also continuous at c


1
Expert's answer
2021-03-10T12:07:56-0500

Solution:

Since "f_n \u2192 f" uniformly on S and each "f_n" is continuous on S, for  given "\u03b5 > 0"

there is a "\delta>0" such that as "|y-x|<\delta" , where "y \in S" , we have

"|f(y)-f(x)|<\frac{\varepsilon}{2}"

For this "\delta>0,\exists" a positive integer "N_{1}" such that as "n \geq N_{1}" , we have

"\left|x_{n}-x\right|<\delta"

Hence, as "n \geq N_{1}" , we have

"\left|f\left(x_{n}\right)-f(x)\right|<\frac{\varepsilon}{2}\ ...(1)"

Also, since "f_{n} \rightarrow f" uniformly on S, given "\varepsilon>0,\exists" a positive integer "N \geq N_{1}" such that as "n \geq N" , we have

"\left|f_{n}(x)-f(x)\right|<\frac{\varepsilon}{2} \text { for all } x \in S"

"\Rightarrow \left|f_{n}\left(x_{n}\right)-f\left(x_{n}\right)\right|<\frac{\varepsilon}{2} \ ... (2)"

By (1) and (2), we obtain that given "\varepsilon>0,\exists" a positive integer N such that as "n \geq N" , we have

"\begin{aligned}

\left|f_{n}\left(x_{n}\right)-f(x)\right| &=\left|f_{n}\left(x_{n}\right)-f\left(x_{n}\right)\right|+\left|f\left(x_{n}\right)-f(x)\right| \\

&<\frac{\varepsilon}{2}+\frac{\varepsilon}{2} \\

&=\varepsilon

\end{aligned}"

"\Rightarrow f_{n}\left(x_{n}\right) \rightarrow f(x)"

Hence, limit function f is also continuous at any point c.

Hence, proved.